# Look Before You Leap: Problem Elaboration Prompting Improves Mathematical Reasoning in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have exhibited impressive performance across NLP tasks. So far they still face challenges in complex reasoning tasks and can be sensitive to input context. Despite significant efforts have been invested in enhancing reasoning process and improving prefix-prompts robustness, the crucial role of problem context has been overlooked. In this study, we propose a new approach to improve the mathematical capacities of LLMs, named Problem Elaboration Prompting (PEP). Specifically, PEP decomposes and elucidates the problem context before reasoning, thus enhancing the global context modeling and reducing the parsing difficulties. Experiments on datasets demonstrate promising performances on complex reasoning and indicate the beneficial impact for ill-formed problems. For instance, with the GPT-3.5 model (text-davinci-003), we observed a 9.93% improvement with greedy decoding and 8.80% improvement with self-consistency on GSM8k compared to the standard CoT. With ChatGPT (turbo) and PEP, we achieve SOTA performances on SVAMP with 86.2% and GSM8k with 90.98%.
[…]
Look Before You Leap: Problem Elaboration Prompting Improves Mathematical Reasoning in Large Language Models
Haoran Liaothanks: Email: liaohaoran@sjtu.edu.cn , Jidong Tian, Shaohua Hu, Hao He, Yaohui Jin
MoE Key Lab of Artificial Intelligence, AI Institute, Shanghai Jiao Tong University
Abstract content
1.   Introduction
@@
Recent large language models (LLMs), such as the GPT-3 model family with 175 billion parameters (Brown et al., 2020, inter alia), have demonstrated remarkable performance across various NLP tasks. Chain-of-thought (CoT) prompting (Wei et al., 2022b; Kojima et al., 2022) successfully elicits reasoning behavior and emergent abilities (Wei et al., 2022a) by explicitly guiding the model to generate intermediate rationales step by step, further promoting the development of artificial general intelligence (AGI). Despite the success, performing multi-hop and compositional reasoning for complex tasks can still face challenges (Hendrycks et al., 2021; Lewkowycz et al., 2022), even when the required knowledge is limited to the scope of primary school (Cobbe et al., 2021).
Figure 1: Possible inputs and outputs of a LLM, including prefix-prompts, problem context, proposed post-problem elaboration, and step-by-step solutions. The complete results and comparisons are presented in Sec. B. In this work, we adopt a zero-shot setting without using exemplars.
Figure 2: An overview of the proposed Problem Elaboration Prompting (PEP) method with standard prompting and CoT prompting. In specific, PEP combs through the problem context before reasoning, aiming to simplify and clarify the given description and asked question to enhance the language model’s comprehension, thereby facilitating a more clear downstream reasoning process.
@@
One area of research aims to improve the quality of reasoning outputs through diverse rationale decoding strategies (Wang et al., 2022c; Suzgun et al., 2022) and iteration-based answer refinement (Saunders et al., 2022; Kim et al., 2023; Zheng et al., 2023). Considering the sensitivity of the model to inputs (Lu et al., 2022b; Wang et al., 2022a; Shi et al., 2023), another area of research focuses on augmenting the robustness of prefix-prompts (Fu et al., 2022; Wang et al., 2022b; Shao et al., 2023). However, the role of problem has been overlooked. Few works, such as Least-to-Most (L2M) prompting, have paid attention to this part (Zhou et al., 2022a; Khot et al., 2022). But L2M only aims to decompose the final question into sub-questions, ignoring the rest problem description.
@@
Recognizing the known conditions and specified intentions of the problem is imperative for subsequent reasoning. The success of previous efforts that have elicited reasoning capabilities in language models, to some extent, benefits from the implicit parsing of problem context (Perez et al., 2020; Zhou et al., 2022a; Wei et al., 2022b). However, most studies assume that the provided information is concise and relevant, while in real-world situations, the input problem could be ill-formed. The language model can be easily distracted by irrelevant context (Shi et al., 2023) and often struggles to capture intricate implied implicatures (Ruis et al., 2022; Chan et al., 2023). Even when the problem is well-formed, it may still be too complex and unsuitable for the language model’s comprehension. For instance, although the language model can be knowledgeable, it may encounter difficulties in identifying what knowledge are required to answer the question (Bian et al., 2023) or correctly integrating intermediate rationales to generate the overall solution (Press et al., 2022).
@@
In this study, we introduce a new method named Problem Elaboration Prompting (PEP), which involves decomposing and elucidating the problem context prior to reasoning. Our method aims to clarify the problem description and enhance the context modeling. We illustrate the possible inputs and outputs for the LLMs in Fig. 1, showing how problem elaboration improves the reasoning of LLMs. PEP adopts a human-like thought process that emphasizes the importance of thoroughly comprehending the problem’s conditions and requirements before engaging in reasoning: look before you leap. PEP can be easily combined with other prompting methods. In addition, our work is also inspired by previous researches on semantic parsing, which suggest parsing the given problem into specific representations, such as Python code or condensed symbols (Gao et al., 2022; Chen et al., 2022; Hu et al., 2023), to facilitate subsequent reasoning.
@@
We present an overview of PEP in Fig. 2, alongside the standard prompting and CoT prompting. We conduct evaluations of the proposed approach on six reasoning datasets and additionally investigate its performance in addressing the distraction problem. PEP consistently outperforms the standard zero-shot baselines (CoT and PoT (Chen et al., 2022)) across datasets, employing both greedy decoding and self-consistency setting. For instance, with GPT-3.5 (text-davinci-003), we observed a 9.93% improvement with greedy decoding on GSM8k (Cobbe et al., 2021), and 8.80% improvement with self-consistency. When using ChatGPT (Turbo), PEP achieves SOTA with 90.98% on GSM8k, and 86.2% on SVAMP (Patel et al., 2021).
@@
    We propose a new method, Problem Elaboration Prompting (PEP), alongside CoT and self-consistency CoT, for improving LLM reasoning abilities.
@@
    We evaluate PEP through extensive experimentation using both text-davinci-003 and turbo, exploring the role of problem context. We hope it could inspire future work.
@@
    We demonstrate the effectiveness of PEP in mitigating the distraction problem, indicating the promising prospect in dealing with ill-formed problems of other types.
2.   Related Works
2.1.   Emergent Abilities and Prompting
@@
With the large-scale unsupervised pre-training technique, LLMs (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023, inter alia) can perform new tasks conditioning on few in-context inputs via mimicking the format of demonstrations (Webson and Pavlick, 2022; Min et al., 2022; Pan et al., 2023). Instruction tuning further improves the generalization of prompts via training a language model to follow the general instructions rather than specific examples (Wei et al., 2021; Ouyang et al., 2022a; Chung et al., 2022).
@@
Instead of directly generating the final answer, chain-of-thought prompting (CoT) methods (Wei et al., 2022b; Creswell et al., 2022; Jung et al., 2022) suggest guiding LLMs to reach the final answer through a step-by-step process, eliciting the emergent reasoning ability (Wei et al., 2022a; Chan et al., 2022; Prystawski and Goodman, 2023). Zelikman et al. (2022) also proposed to use prompts to augment the trainset with rationales. However, such methods typically necessitate an LLM larger than 100B (Wei et al., 2022b), making it difficult to directly apply on small models (Zhang et al., 2023).
2.2.   Improving CoT Reasoning
@@
Various techniques have been proposed to improve the standard CoT (Wang et al., 2022c; Suzgun et al., 2022; Saunders et al., 2022; Gao et al., 2022; Kim et al., 2023; Yao et al., 2023), most with a particular focus on controlling the decoding process and validating intermediate answers. Significant efforts have also been made to improve the prompts selection and ensemble (Fu et al., 2022; Lu et al., 2022a; Zhou et al., 2022b; Shao et al., 2023), to enhance the generation robustness. Besides, increasing the diversity of prompts, questions, or rationales shows notable advantages for the reasoning (Zhang et al., 2022b; Wang et al., 2022b).
Our work is similar to the study conducted by Lampinen et al. (2022), which explores the impact of explanations in context. However, it investigates the post-answer explanation about rationales, while we focus on elaborating the problem context.
2.3.   Collaborating CoT with agents.
Recent works attempt to introduce the concept of agents to enhance the reasoning, explicitly prompting the LLM with external tools (Gao et al., 2022) to play different roles in undertaking decomposed tasks  (Shinn et al., 2023; Park et al., 2023; Wang et al., 2023a). It typically includes planning, cooperation, competition, and interaction with the environment (Xi et al., 2023). Exploring the decision-making process (Jung et al., 2022) with specific agents has also shown promising improvements, including tree-structure (Yao et al., 2023) and graph-structure (McNichols et al., 2023; Cao, 2023).
Our work can also be regarded as a type of LLM-based agent system. Although existing approaches also have the concept of analysis, they focus more on observing the state of the environment and reflecting the feedback (Zheng et al., 2023; Qian et al., 2023; Wu et al., 2023; Miao et al., 2023), while PEP focuses on analyzing and parsing the problem.
3.   Problem Elaboration Prompting
@@
A beneficial cognitive paradigm of humanity is the ability of thoughtful contemplation, as exemplified by a common figurative: look before you leap. In this work, we propose to simulate such a thinking process by prompting the model to elaborate on the problem before generating rationales and answers. In other words, it is crucial for the language model to ensure it has a thorough understanding of the problem space, encompassing the given conditions and the asked question, avoiding a hasty generation and exploration within the answer space (Table 1).
@@
A: Three times as many video games as Brian has is 3*20=<<3*20=60>>60 video games. Five fewer than 3 times as many video games as Brian has is 60-5=<<60-5=55>>55 video games.
Final answer: 55.
(Wrong: ignore the precondition placed at end.)
@@
(+Elaboration; Declaratives):
Segment 1: Bobby has 5 fewer than 3 times as many video games as Brian does. Rephrased: Bobby’s number of video games is 3 times Brian’s number minus 5.
Equation: B = 3B - 5
Segment 2: Brian has 20 video games.
Equation: B = 20
Segment 3: Brian lost 5 video games right before the comparison was made.
Equation: B = 20 - 5
@@
Final answer: 40.
(Correct: process conditions in suitable order.)
Table 1: An example of PEP with turbo. After capturing the relevant conditions presented upfront, the vanilla CoT wrongly engages in hasty reasoning (red), neglecting the important prerequisite conditions placed afterwards (cyan). Meanwhile, PEP benefits from the preprocessing elaboration and composites conditions in the correct order (blue).
3.1.   Method Formulation
@@
In the PEP-aid language model, we suggest to pre-process the given problem $P$ by decomposing and elucidating the context into smaller and more concise segments to enhance LLMs’ comprehension: $P^{\prime}=(\{x_{1}^{\prime},x_{2}^{\prime},\cdots,x_{m}^{\prime}\},y_{0}^{\prime})=f_{M}(X,y_{0};p_{e})$, where $m\geq i$ and $p_{e}$ is a specific instruction. Then, the LLM can continue its reasoning by $y_{i}=f_{M}(X,y_{0};p_{t},\hat{P}_{k}|P^{\prime},y_{<i})$ utill reaching $y_{t}$. Thus, PEP can easily be combined with previous prompting methods.
3.2.   Designing Principles
@@
The concept of decomposition is widely spread in many previous works. For example, the L2M prompting (Zhou et al., 2022a) suggests breaking down the asked questions into sub-questions and then prompts the model to solve them one by one. The Plan-and-Solve prompting (Wang et al., 2023b) suggests generating an overall plan before reasoning. In contrast, PEP takes a different approach by breaking down the entire problem into simpler segments, rather than just decomposing the final question. Furthermore, PEP does not require deducing a plan or inferring the sub-questions. Instead, it focuses on organizing and clarifying the existing information from the problem.
Meanwhile, it could be beneficial to elucidate the segments following decomposition, as it elicits the model to organize existing information in a comprehensive view of the problem and recognize the underlying implicatures of the question. Moreover, it introduces diversity into the context, which has been demonstrated to enhance reasoning (Zhang et al., 2022b; Wang et al., 2022b), thereby mitigating the risk of relying on specific words or descriptions as shortcuts. A break-down analysis of our designed principles is presented in Sec. 4.4.
Q/A pair Declaratives Interrogative Mixed
19.5 47.0 23.0 10.5
Table 2: Distribution of PEP forms. We annotated the corresponding types in respective cases.
3.3.   Instruction Generation
We utilize the zero-shot setting to prompt the language models since the elaboration of different problems can be diverse. We apply LLM to generate alternatives. Then, the instructions are evaluated on 200 points randomly sampled from the GSM8k trainset. The final selected instruction is “Decompose the given question into smaller segments, elucidating each segment as you rephrase it.” The details of the selection process and all used prompts are provided in the Appendix.
To investigate the behavior of PEP, we randomly selected 200 instances from the experiments with a manual categorization. PEP primarily utilizes question-answer pairs, declarative sentences, and interrogative sentences to review and examine the problem context (see Table 2). There are around 10.5% of instances that can be mixed with sub-questions, planning instructions, or intertwined rationales. We find it primarily occurs when the questions themselves contain instructions or options.
4.   Experiment
4.1.   Setup
Dataset #Test #AT domain
SingleEq 508 free math
GSM8k 1319 free math
AQuA 254 options math
SVAMP 1000 free math
MMLU-M 500 options math
GSMIC-1k 1000 free distraction
Table 3: Statistical information of used datasets. #Test refers to the number of test samples and #AT denotes the answer types.

### v2

@@ Abstract
Large language models (LLMs) still grapple with complex tasks like mathematical reasoning. Despite significant efforts invested in improving prefix prompts or reasoning process, the crucial role of problem context might have been neglected. Accurate recognition of inputs is fundamental for solving mathematical tasks, as ill-formed problems could potentially mislead LLM’s reasoning. In this study, we propose a new approach named Problem Elaboration Prompting (PEP) to enhance the mathematical capacities of LLMs. Specifically, PEP decomposes and elucidates the problem context before reasoning, therefore enhancing the context modeling and parsing efficiency. Experiments across datasets and models demonstrate promising performances: (1) PEP demonstrates an overall enhancement in various mathematical tasks. For instance, with the GPT-3.5 model, PEP exhibits improvements of 9.93% and 8.80% on GSM8k through greedy decoding and self-consistency, respectively. (2) PEP can be easily implemented and integrated with other prompting methods. (3) PEP shows particular strength in handling distraction problems (example in Fig. 1).
Figure 1: We proposed Problem Elaboration Prompting (PEP) for enhancing problem context, thereby improving subsequent reasoning. As depicted in the example, PEP decouples spurious relationships and refines statements, preventing downstream distraction errors.
1 Introduction
@@
Recent large language models (LLMs), such as the GPT-3 model family with 175 billion parameters (Brown et al., 2020, inter alia), have demonstrated remarkable performance across various NLP tasks. Chain-of-thought (CoT) prompting (Wei et al., 2022b; Kojima et al., 2022) successfully elicits reasoning behavior and emergent abilities (Wei et al., 2022a) by explicitly guiding the model to generate intermediate rationales step by step, further promoting the development of artificial general intelligence (AGI). Despite the success, performing multi-hop and compositional reasoning for complex tasks like mathematical solving can still face challenges (Hendrycks et al., 2021; Lewkowycz et al., 2022), even when the required knowledge is limited to the scope of primary school (Cobbe et al., 2021).
Figure 2: An overview of the proposed PEP and other problem-related methods. Rather than creating sub-questions or plans to guide subsequent reasoning, PEP focuses on clarifying and enriching the problem context, i.e., PEP can be integrated with these methods.
@@
One area of research aims to improve the quality of reasoning outputs through diverse rationale decoding strategies (Wang et al., 2022c; Suzgun et al., 2022) and iteration-based answer refinement (Saunders et al., 2022; Kim et al., 2023; Zheng et al., 2023). Considering the sensitivity of the model to inputs (Lu et al., 2022b; Wang et al., 2022a; Shi et al., 2023), another area of research focuses on augmenting the robustness of prefix-prompts (Fu et al., 2022; Wang et al., 2022b; Shao et al., 2023). However, the role of problem has been overlooked.
@@
Most studies assume that the provided information is concise and relevant, while in real-world situations, the inputs could be ill-formed. For instance, LMs can be easily distracted by irrelevant context (Shi et al., 2023) and often struggle to capture intricate implied implicatures (Ruis et al., 2022; Chan et al., 2023). Besides, even when the problem is well-formed, it may still be complex and unsuitable for the LM’s comprehension. For instance, although the language model can be knowledgeable, it may encounter difficulties in identifying what knowledge are required to answer the question (Bian et al., 2023) or correctly integrating intermediate rationales to generate the overall solution (Press et al., 2022).
Several works have noticed the crucial role of problem, suggesting pre-processing before reasoning. Least-to-Most (L2M) (Zhou et al., 2022a) proposes to decompose the final asked question into simpler sub-questions, Plan-and-Solve (PaS)  (Wang et al., 2023b) requires a preliminary global plan, while Self-ask (Press et al., 2022) suggests a dynamic asking strategy before each step of generating rationales. However, these methods primarily focus on decomposing questions or creating guidance, without understanding or discernment of the problem itself. Therefore, they could also potentially be misled by ill-formed problems.
@@
In this study, we introduce a new method named Problem Elaboration Prompting (PEP), which involves decomposing and elucidating the problem context prior to reasoning. Our method aims to clarify the problem description and enhance the context modeling, rather than creating specific guidance. We illustrated an overview of PEP in Fig 2, along with a comparison with other problem-related prompting methods. Specifically, PEP adopts a human-like thought process that emphasizes the importance of thoroughly comprehending the problem’s conditions and requirements before engaging in reasoning: look before you leap. PEP is also inspired by previous researches on semantic parsing, which suggest parsing the given problem into specific representations, such as Python code or condensed symbols (Gao et al., 2022; Chen et al., 2022; Hu et al., 2023), to facilitate subsequent reasoning.
@@
We conduct evaluations of the proposed approach on four mathematical datasets and additionally investigate its performance in addressing the distraction problems (Shi et al., 2023). Both zero-shot and few-shot PEPs demonstrate an overall enhancement across datasets, models and answer types, employing greedy decoding and self-consistency settings. For instance, with GPT-3.5, we observed an improvement of 9.93% and 8.80% on GSM8k (Cobbe et al., 2021) with greedy decoding and self-consistency, separately. When using ChatGPT, PEP achieves SOTAs with 98.23% on SingleEq (Koncel-Kedziorski et al., 2015), and 88.7% on SVAMP (Patel et al., 2021).
@@
    We propose a new method, Problem Elaboration Prompting (PEP), for enhancing LLM’s reasoning. It is easy to implement and integrate other prompting methods.
@@
    We evaluate PEP through extensive experiments using both open-source LMs and GPT model family, exploring the role of problem context for reasoning.
@@
    We demonstrate PEP is effective in mitigating the distraction problem, indicating the promising prospect in dealing with ill-formed problems of other types.
2 Related Works
2.1 Emergent Abilities and Prompting
@@
With the large-scale unsupervised pre-training technique, LLMs (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a, inter alia) can perform new tasks conditioning on few in-context inputs via mimicking the format of demonstrations (Webson & Pavlick, 2022; Min et al., 2022; Pan et al., 2023). Instruction tuning further improves the generalization of prompts via training a language model to follow the general instructions rather than specific examples (Wei et al., 2021; Ouyang et al., 2022a; Chung et al., 2022).
@@
Instead of directly generating the final answer, chain-of-thought prompting (CoT) methods (Wei et al., 2022b; Creswell et al., 2022; Jung et al., 2022) suggest guiding LLMs to reach the final answer through a step-by-step process, eliciting the emergent reasoning ability (Wei et al., 2022a; Chan et al., 2022; Prystawski & Goodman, 2023). Zelikman et al. (2022) also proposed to use prompts to augment the trainset with rationales. However, such methods typically necessitate an LLM larger than 100B (Wei et al., 2022b), making it difficult to directly apply on small models (Zhang et al., 2023).
2.2 Improving CoT Reasoning
@@
Various techniques have been proposed to improve the standard CoT (Wang et al., 2022c; Suzgun et al., 2022; Saunders et al., 2022; Gao et al., 2022; Kim et al., 2023; Yao et al., 2023), most with a particular focus on controlling the decoding process and validating intermediate answers. Significant efforts have also been made to improve the prompts selection and ensemble (Fu et al., 2022; Lu et al., 2022a; Zhou et al., 2022b; Shao et al., 2023), to enhance the generation robustness. Besides, increasing the diversity of prompts, questions, or rationales shows notable advantages for the reasoning (Zhang et al., 2022; Wang et al., 2022b).
PEP shares similarity with the study conducted by Lampinen et al. (2022), which explores the impact of explanations in context. However, it focus on the post-answer explanation, while we focus on elaborating the problem. PEP can also be regarded as a type of LLM-based agent system, sharing the concepts of analysis and decomposition (Shinn et al., 2023; Park et al., 2023; Wang et al., 2023a). While they primarily focus on observing environment and reflecting the feedback (Zheng et al., 2023; Qian et al., 2023; Wu et al., 2023; Miao et al., 2023), PEP focuses on comprehending and parsing the context.
3 Problem Elaboration Prompting
@@
A beneficial cognitive paradigm of humanity is the ability of thoughtful contemplation, as exemplified by a common figurative: look before you leap. In this work, we propose to simulate such a thinking process by prompting the model to elaborate on the problem before generating or reasoning. It is crucial for the LLM to ensure it has a thorough understanding of the problem space, encompassing the given conditions and the asked question, avoiding a hasty generation and exploration within the answer space (Table 1).
@@
A: Three times as many video games as Brian has is 3*20=60 video games. Five fewer than 3 times as many video games as Brian has is 60-5=55 video games.
Final answer: 55. (Wrong: ignore the precondition placed at end.)
@@
(+Elaboration):
Segment 1: Bobby has 5 fewer than 3 times as many video games as Brian does. Rephrased: Bobby’s number of video games is 3 times Brian’s number minus 5. Equation: B = 3B - 5
Segment 2: Brian has 20 video games. Equation: B = 20
Segment 3: Brian lost 5 video games right before the comparison was made. Equation: B = 20 - 5
@@
Final answer: 40. (Correct: process conditions in suitable order.)
Table 1: Recognizing the problem is imperative for reasoning. As illustrated, CoT wrongly engages in hasty reasoning (red), neglecting the prerequisite conditions placed afterwards (cyan). Meanwhile, PEP utilizes conditions in the correct order (blue).
3.1 Method Formulation
@@
In the PEP-aid language model, we suggest to pre-process the given problem $P$ by decomposing and elucidating the context into smaller and more concise segments to enhance LLMs’ comprehension: $P^{\prime}=(\{x_{1}^{\prime},x_{2}^{\prime},\cdots,x_{m}^{\prime}\},y_{0}^{\prime})=f_{M}(X,y_{0};p_{e})$, where $m\geq i$ and $p_{e}$ is a specific instruction. Then, the LLM can continue its reasoning by $y_{i}=f_{M}(X,y_{0};p_{t},\hat{P}_{k}|P^{\prime},y_{<i})$ utill reaching $y_{t}$. Thus, PEP can be easily combined with previous prompting methods.
3.2 Designing Principles
@@
The concept of decomposition is widely spread in many previous works. Except for introduced problem-related methods, recent community adopts a decomposition approach in different fields to apply LMs to solve complex problems (Wei et al., 2022b; Khot et al., 2022; Liang et al., 2023; Hong et al., 2023), bridging the compositionality gap of powerful LMs (Press et al., 2022). In contrast, PEP takes a different approach by breaking down the entire problem into simpler segments, rather than creating sub-questions or decomposing the reasoning process. It focuses on organizing and clarifying the existing information from the problem.
Meanwhile, it could be beneficial to elucidate the segments following decomposition, as it elicits the model to organize existing information in a comprehensive view of the problem and recognize the underlying implicatures of the question. Moreover, it introduces diversity into the context, which has been demonstrated to enhance reasoning (Zhang et al., 2022; Wang et al., 2022b), thereby mitigating the risk of relying on specific words or descriptions as shortcuts. A break-down analysis of our designed principles is presented in Sec. 4.4.
3.3 Prompts Generation
Since the elaboration can be diverse, We evaluate and select the instruction of zero-shot PEP based on a subset of GSM8k-train of 200 points (see Sec A). The selected instruction is “Decompose the given question into smaller segments, elucidating each segment as you rephrase it.” We further adopt the exemplars from Zhou et al. (2022a) and adjusted them for different methods for fairness. All instructions and exemplars can be found in the appendix B.
To investigate the behavior of PEP, we randomly selected 200 instances from the experiments with a manual categorization. PEP primarily utilizes question-answer pairs, declarative sentences, and interrogative sentences to review and examine the problem context. There are around 10.5% of instances that can be mixed with sub-questions, planning instructions, or intertwined rationales. We find it primarily occurs when the questions themselves contain instructions or options.
4 Experiment
4.1 Setup
Dataset Num Answer
SingleEq 508 free
GSM8k 1319 free
AQuA 254 options
SVAMP 1000 free
GSMIC-1k 1000 free
Table 2: Datasets Stats.